# 3DIS: Depth-Driven Decoupled Instance Synthesis for Text-to-Image Generation

### v1

@@ 1 Introduction
With the rapid advancement of text-to-image generation technologies, there is a growing interest in achieving more controllable outputs, which are now widely utilized in artistic creation (Wang et al., 2024; Li et al., 2024a): (i) Image-conditional generation techniques, e.g., ControlNet (Zhang et al., 2023), allow users to generate images based on inputs like depth maps or sketches. (ii) Multi-instance generation (MIG) methods, e.g., GLIGEN (Li et al., 2023b) and MIGC (Zhou et al., 2024a), enable users to define layouts and detailed attributes for each instance within the generated images.
However, despite the importance of MIG in controllable generation, these methods have not been widely adopted across popular foundational models like SD2 (Rombach et al., 2023) and SDXL (Podell et al., 2023), unlike the more widely integrated ControlNet. Current state-of-the-art MIG methods mainly rely on the less capable SD1.5 (Rombach et al., 2022) model.
@@
The 3DIS architecture comprises three key components: (i) Scene Depth Map Generation: We developed the first layout-controllable text-to-depth generation model by integrating a well-designed adapter into LDM3D (Stan et al., 2023). This integration facilitates the generation of precise, depth-informed layouts based on instance conditions. (ii) Layout Control: We introduce a method to leverage pretrained ControlNet for seamless integration of the generated scene depth map into the generative process. By filtering out high-frequency information from ControlNet’s feature maps, we enhance the integration of low-frequency global scene semantics, thereby improving the coherence and visual appeal of the generated images. (iii) Detail Rendering: Our method performs Cross-Attention operations separately for each instance to achieve precise rendering of specific attributes (e.g., category, color, texture) while avoiding attribute leakage. Additionally, we use SAM for semantic segmentation on the scene depth map, optimizing instance localization and resolving conflicts from overlapping bounding boxes. This advanced approach significantly improves the rendering of detailed and accurate multi-instance images.
We conducted extensive experiments on two benchmarks to evaluate the performance of 3DIS: (i) COCO-Position (Lin et al., 2015; Zhou et al., 2024a): Evaluated the layout accuracy and coarse-grained category attributes of the scene depth maps. (ii) COCO-MIG (Zhou et al., 2024a): Assessed the fine-grained rendering capabilities. The results indicate that 3DIS excels in creating superior scenes while preserving the accuracy of fine-grained attributes during detailed rendering. On the COCO-Position benchmark, 3DIS achieved a 16.3% improvement in AP75 compared to the previous state-of-the-art method, MIGC. On the COCO-MIG benchmark, our training-free detail rendering approach improved the Instance Attribute Success Ratio by 35% over the training-free method Multi-Diffusion (Bar-Tal et al., 2023) and by 5.5% over the adapter-based method InstanceDiffusion (Wang et al., 2024). Furthermore, the 3DIS framework can be seamlessly integrated with off-the-shelf adapters like GLIGEN and MIGC, thereby enhancing their rendering capabilities.
@@ 2 Related Work
Controllable Text-to-Image Generation. With the rapid advancements in diffusion models (Zhou et al., 2023; Lu et al., 2024; Zhao et al., 2024a; b; Xie et al., 2024), text-to-image technologies have reached a level where they can produce high-quality images (Rombach et al., 2022; 2023; Podell et al., 2023; Lu et al., 2023). Researchers are now increasingly focused on enhancing their control over the generated content. Numerous approaches have been developed to improve this control. ControlNet (Zhang et al., 2023) incorporates user inputs such as depth maps and edge maps by training an additional side network, allowing for precise layout control in image generation. Methods like IPAdapter (Ye et al., 2023) and PhotoMaker (Li et al., 2024b) generate corresponding images based on user-provided portraits. Techniques such as ELITE (Wei et al., 2023) and SSR-Encoder (Zhang et al., 2024) enable networks to accept specific conceptual image inputs for better customization. Additionally, MIGC (Zhou et al., 2024a; b) and InstanceDiffusion (Wang et al., 2024) allow networks to generate images based on user-specified layouts and instance attribute descriptions, defining this task as Multi-Instance Generation (MIG), which is the focal point of this paper.
Multi-Instance Generation (MIG). MIG involves generating each instance based on a given layout and detailed attribute descriptions, while maintaining overall image harmony. Current MIG methods primarily use Stable Diffusion (SD) architectures, classified into three categories: 1) Training-free methods: Techniques like BoxDiffusion (Xie et al., 2023) and RB (Xiao et al., 2023) apply energy functions to attention maps, enabling zero-shot layout control by converting spatial guidance into gradient inputs. Similarly, Multi-Diffusion (Bar-Tal et al., 2023) generates instances separately and then combines them according to user-defined spatial cues, enhancing control over orientation and arrangement. 2) Adapter methods: Approaches like GLIGEN (Li et al., 2023b) and InstanceDiffusion (Wang et al., 2024) integrate trainable gated self-attention layers into the U-Net (Ronneberger et al., 2015), improving layout assimilation and instance fidelity. MIGC (Zhou et al., 2024a) further divides the task, using an enhanced attention mechanism to generate each instance precisely before integration. 3) SD-tuning methods: Reco (Yang et al., 2023) and Ranni (Feng et al., 2024) add instance position data to text inputs and fine-tune both CLIP and U-Net, allowing the network to utilize positional cues for more precise image synthesis. Previous methods entangled instance positioning with attribute rendering, complicating the training of a robust instance renderer. Our approach decouples this process into adapter-controlled scene depth map generation and training-free detail rendering. This separation allows the adapter to only handle instance positioning and coarse attributes, while leveraging the generative priors of pre-trained models, enhancing both flexibility and performance.
@@ 3.3 Scene Depth Map Generation
Choosing the text-to-depth model. Upon investigation, we identified RichDreamer (Qiu et al., 2024) and LDM3D (Stan et al., 2023) as the primary models for text-to-depth generation. RichDreamer fine-tunes the pretrained RGB Stable Diffusion (SD) model to generate 3D information, specifically depth and normal maps, while LDM3D enables SD to produce both RGB images and depth maps simultaneously. Experimental comparisons show LDM3D outperforms RichDreamer in complex scenes, likely due to its concurrent RGB and depth map generation. This dual capability preserves RGB image quality while enhancing depth map generation, making LDM3D our preferred model for text-to-depth generation.
Fine-tuning the text-to-depth model. In contrast to RGB images, depth maps typically prioritize the restoration of low-frequency components over high-frequency details. For instance, while a texture-rich skirt requires intricate details for RGB image generation, its corresponding depth map remains relatively smooth. Therefore, we aim to enhance the model’s ability to recover low-frequency content. Low-frequency components often indicate significant redundancy among adjacent pixels. To simulate this characteristic, we implemented an augmented pyramid noise strategy (Kasiopy, 2023), which involves downsampling and then upsampling randomly sampled noise $\epsilon$ to create patterns with high redundancy between adjacent pixels. We used the original SD training loss (Rombach et al., 2022) to fine-tune our text-to-depth model $\theta$, but adjusted the model to predict this patterned noise $\epsilon_{\text{pyramid}}$ with the text prompt $\bm{c}$:
@@
Training the Layout-to-depth adapter. Similar to previous methodologies (Zhou et al., 2024a; Li et al., 2023b; Wang et al., 2024), we incorporated an adapter into our fine-tuned text-to-depth model, enabling layout-to-depth generation, specifically leveraging the state-of-the-art MIGC (Zhou et al., 2024a) model. Unlike earlier approaches, our method for generating depth maps does not rely on detailed descriptions of specific instance attributes, such as material or color. Consequently, we have augmented the dataset used for MIGC by eliminating fine-grained attribute descriptions from the instance data, thus focusing more on the structural properties of individual instances and the overall scene composition. The training process for the adapter $\theta^{\prime}$ can be expressed as:
@@ 3.4 Global Scene Control
Injecting depth maps with ControlNet. After generating scene depth maps with our layout-to-depth models, we employed the widely adopted ControlNet (Zhang et al., 2023) model to incorporate global scene information. Scene depth maps focus on overall scene structure, without requiring fine-grained detail. Thus, although the base model produces 512x512 resolution maps, they can be upsampled to 768x768, 1024x1024, or higher (see Fig. 3 and Fig. 4, e.g., SD2 and SDXL). Since most generative models have depth ControlNet versions, these maps can be applied across various models, ensuring accurate instance placement and mitigating omission issues.
@@ 3.5 Details rendering
SAM-Enhancing Instance Location. While mering rendering results, acquiring precise instance locations helps prevent attribute leakage between overlapping bounding boxes and maintains structural consistency with the instances in the scene depth maps. Consequently, we employ the SAM (Kirillov et al., 2023) model to ascertain the exact position of each instance. For an instance $\bm{i}$, by utilizing our generated scene depth map $\mathbf{m}_{scene}$ alongside its corresponding bounding box $\bm{b}_{i}$, we can segment the specific shape mask $\mathbf{m}_{i}$ of this instance, thereby facilitating subsequent merging:
@@
Merging rendering results. We employ the precise mask $\mathbf{m}_{i}$ obtained from SAM to constrain the rendering results of instance $\bm{i}$ to its own region, ensuring no influence on other instances. Specifically, we construct a new mask $\mathbf{m}_{i}^{\prime}$ by assigning a value of $\alpha$ to the areas where $\mathbf{m}_{i}$ equals $1$, while setting all other regions to $\infty$- ∞. Simultaneously, we assign a background value of $\beta$ to the global rendering $\mathbf{r}_{c}$ through a mask $\mathbf{m}_{c}^{\prime}$. By applying the softmax function to the set $\{\mathbf{m}_{1}^{\prime},\mathbf{m}_{2}^{\prime},\ldots,\mathbf{m}_{n}^{\prime},\mathbf{m}_{c}^{\prime}\}$, we derive the spatial weights $\{\mathbf{m}_{1}^{\prime\prime},\mathbf{m}_{2}^{\prime\prime},\ldots,\mathbf{m}_{n}^{\prime\prime},\mathbf{m}_{c}^{\prime\prime}\}$ for each rendering instance. At each Cross Attention layer, the output can be expressed as follows to render the whole scene:
@@ 4.1 Implement Details
Tuning of text-to-depth models. We utilized a training set comprising 5,878 images from the LAION-art dataset (Schuhmann et al., 2021), selecting only those with a resolution exceeding 512x512 pixels and an aesthetic score of $\geq{8.0}$. Depth maps for each image were generated using Depth Anything V2 (Yang et al., 2024). Given the substantial noise present in the text descriptions associated with the images in LAION-art, we chose to produce corresponding image captions using BLIP2 (Li et al., 2023a). We employed pyramid noise (Kasiopy, 2023) to fine-tune the LDM3D model for 2,000 steps, utilizing the AdamW (Kingma & Ba, 2017) optimizer with a constant learning rate of $1e^{-4}$, a weight decay of $1e^{-2}$, and a batch size of 320.
Training of the layout-to-depth adapter. We adopted the MIGC (Zhou et al., 2024a) architecture as the adapter for layout control. In alignment with this approach, we utilized the COCO dataset (Lin et al., 2015) for training. We employed Stanza (Qi et al., 2020) to extract each instance description from the corresponding text for every image and used Grounding-DINO (Liu et al., 2023) to obtain the image layout. Furthermore, we augmented each instance’s description by incorporating modified versions that omitted adjectives, allowing our layout-to-depth adapter to focus more on global scene construction and the coarse-grained categories and structural properties of instances. We maintain the same batch size, learning rate, and other parameters as the previous work.
@@ 4.2 Experiment Setup
Baselines. We compared our proposed 3DIS method with state-of-the-art Multi-Instance Generation approaches. The methods involved in the comparison include training-free methods: BoxDiffusion (Xie et al., 2023) and MultiDiffusion (Bar-Tal et al., 2023); and adapter-based methods: GLIGEN (Li et al., 2023b), InstanceDiffusion (Wang et al., 2024), and MIGC (Zhou et al., 2024a).
Evaluation Benchmarks. We conducted experiments using two widely adopted benchmarks, COCO-position (Lin et al., 2015) and COCO-MIG (Zhou et al., 2024a), to assess the performance of models in different aspects of instance generation. The COCO-position benchmark emphasizes the evaluation of a model’s capacity to control the spatial arrangement of instances, as well as their high-level categorical attributes. In contrast, the COCO-MIG benchmark is designed to test a model’s ability to precisely render fine-grained attributes for each generated instance. To rigorously compare the models’ performance in handling complex scene layouts, we concentrated our analysis on the COCO-position benchmark, specifically focusing on layouts containing five or more instances. For a comprehensive evaluation, each model generated 750 images across both benchmarks.

### v2

@@ 1 Introduction
With the rapid advancement of text-to-image generation technologies, there is a growing interest in achieving more controllable outputs, which are now widely utilized in artistic creation (instancediffusion; li2024controlnet++): (i) Image-conditional generation techniques, e.g., ControlNet (controlnet), allow users to generate images based on inputs like depth maps or sketches. (ii) Multi-instance generation (MIG) methods, e.g., GLIGEN (gligen) and MIGC (migc), enable users to define layouts and detailed attributes for each instance within the generated images.
However, despite the importance of MIG in controllable generation, these methods have not been widely adopted across popular foundational models like SD2 (SD2) and SDXL (podell2023sdxl), unlike the more widely integrated ControlNet. Current state-of-the-art MIG methods mainly rely on the less capable SD1.5 (stablediffusion) model.
@@
The 3DIS architecture comprises three key components: (i) Scene Depth Map Generation: We developed the first layout-controllable text-to-depth generation model by integrating a well-designed adapter into LDM3D (stan2023ldm3d). This integration facilitates the generation of precise, depth-informed layouts based on instance conditions. (ii) Layout Control: We introduce a method to leverage pretrained ControlNet for seamless integration of the generated scene depth map into the generative process. By filtering out high-frequency information from ControlNet’s feature maps, we enhance the integration of low-frequency global scene semantics, thereby improving the coherence and visual appeal of the generated images. (iii) Detail Rendering: Our method performs Cross-Attention operations separately for each instance to achieve precise rendering of specific attributes (e.g., category, color, texture) while avoiding attribute leakage. Additionally, we use SAM for semantic segmentation on the scene depth map, optimizing instance localization and resolving conflicts from overlapping bounding boxes. This advanced approach significantly improves the rendering of detailed and accurate multi-instance images.
We conducted extensive experiments on two benchmarks to evaluate the performance of 3DIS: (i) COCO-Position (coco; migc): Evaluated the layout accuracy and coarse-grained category attributes of the scene depth maps. (ii) COCO-MIG (migc): Assessed the fine-grained rendering capabilities. The results indicate that 3DIS excels in creating superior scenes while preserving the accuracy of fine-grained attributes during detailed rendering. On the COCO-Position benchmark, 3DIS achieved a 16.3% improvement in AP75 compared to the previous state-of-the-art method, MIGC. On the COCO-MIG benchmark, our training-free detail rendering approach improved the Instance Attribute Success Ratio by 35% over the training-free method Multi-Diffusion (multidiffusion) and by 5.5% over the adapter-based method InstanceDiffusion (instancediffusion). Furthermore, the 3DIS framework can be seamlessly integrated with off-the-shelf adapters like GLIGEN and MIGC, thereby enhancing their rendering capabilities.
@@ 2 Related Work
Controllable Text-to-Image Generation. With the rapid advancements in diffusion models (pydiff; lu2024mace; zhao2023wavelet; zhao2024learning; Xie2024AddSRAD), text-to-image technologies have reached a level where they can produce high-quality images (stablediffusion; SD2; podell2023sdxl; lu2023tf). Researchers are now increasingly focused on enhancing their control over the generated content. Numerous approaches have been developed to improve this control. ControlNet (controlnet) incorporates user inputs such as depth maps and edge maps by training an additional side network, allowing for precise layout control in image generation. Methods like IPAdapter (ip-adapter) and PhotoMaker (li2024photomaker) generate corresponding images based on user-provided portraits. Techniques such as ELITE (elite) and SSR-Encoder (ssr) enable networks to accept specific conceptual image inputs for better customization. Additionally, MIGC (migc; migc++) and InstanceDiffusion (instancediffusion) allow networks to generate images based on user-specified layouts and instance attribute descriptions, defining this task as Multi-Instance Generation (MIG), which is the focal point of this paper.
Multi-Instance Generation (MIG). MIG involves generating each instance based on a given layout and detailed attribute descriptions, while maintaining overall image harmony. Current MIG methods primarily use Stable Diffusion (SD) architectures, classified into three categories: 1) Training-free methods: Techniques like BoxDiffusion (boxdiff) and RB (rb) apply energy functions to attention maps, enabling zero-shot layout control by converting spatial guidance into gradient inputs. Similarly, Multi-Diffusion (multidiffusion) generates instances separately and then combines them according to user-defined spatial cues, enhancing control over orientation and arrangement. 2) Adapter methods: Approaches like GLIGEN (gligen) and InstanceDiffusion (instancediffusion) integrate trainable gated self-attention layers into the U-Net (Unet), improving layout assimilation and instance fidelity. MIGC (migc) further divides the task, using an enhanced attention mechanism to generate each instance precisely before integration. 3) SD-tuning methods: Reco (reco) and Ranni (ranni) add instance position data to text inputs and fine-tune both CLIP and U-Net, allowing the network to utilize positional cues for more precise image synthesis. Previous methods entangled instance positioning with attribute rendering, complicating the training of a robust instance renderer. Our approach decouples this process into adapter-controlled scene depth map generation and training-free detail rendering. This separation allows the adapter to only handle instance positioning and coarse attributes, while leveraging the generative priors of pre-trained models, enhancing both flexibility and performance.
@@ 3.3 Scene Depth Map Generation
Choosing the text-to-depth model. Upon investigation, we identified RichDreamer (richdreamer) and LDM3D (stan2023ldm3d) as the primary models for text-to-depth generation. RichDreamer fine-tunes the pretrained RGB Stable Diffusion (SD) model to generate 3D information, specifically depth and normal maps, while LDM3D enables SD to produce both RGB images and depth maps simultaneously. Experimental comparisons show LDM3D outperforms RichDreamer in complex scenes, likely due to its concurrent RGB and depth map generation. This dual capability preserves RGB image quality while enhancing depth map generation, making LDM3D our preferred model for text-to-depth generation.
Fine-tuning the text-to-depth model. In contrast to RGB images, depth maps typically prioritize the restoration of low-frequency components over high-frequency details. For instance, while a texture-rich skirt requires intricate details for RGB image generation, its corresponding depth map remains relatively smooth. Therefore, we aim to enhance the model’s ability to recover low-frequency content. Low-frequency components often indicate significant redundancy among adjacent pixels. To simulate this characteristic, we implemented an augmented pyramid noise strategy (pyramidnoise), which involves downsampling and then upsampling randomly sampled noise $\epsilon$ to create patterns with high redundancy between adjacent pixels. We used the original SD training loss (stablediffusion) to fine-tune our text-to-depth model $\theta$, but adjusted the model to predict this patterned noise $\epsilon_{\text{pyramid}}$ with the text prompt $\bm{c}$:
@@
Training the Layout-to-depth adapter. Similar to previous methodologies (migc; gligen; instancediffusion), we incorporated an adapter into our fine-tuned text-to-depth model, enabling layout-to-depth generation, specifically leveraging the state-of-the-art MIGC (migc) model. Unlike earlier approaches, our method for generating depth maps does not rely on detailed descriptions of specific instance attributes, such as material or color. Consequently, we have augmented the dataset used for MIGC by eliminating fine-grained attribute descriptions from the instance data, thus focusing more on the structural properties of individual instances and the overall scene composition. The training process for the adapter $\theta^{\prime}$ can be expressed as:
@@ 3.4 Global Scene Control
Injecting depth maps with ControlNet. After generating scene depth maps with our layout-to-depth models, we employed the widely adopted ControlNet (controlnet) model to incorporate global scene information. Scene depth maps focus on overall scene structure, without requiring fine-grained detail. Thus, although the base model produces 512x512 resolution maps, they can be upsampled to 768x768, 1024x1024, or higher (see Fig. 3 and Fig. 4, e.g., SD2 and SDXL). Since most generative models have depth ControlNet versions, these maps can be applied across various models, ensuring accurate instance placement and mitigating omission issues.
@@ 3.5 Details rendering
SAM-Enhancing Instance Location. While mering rendering results, acquiring precise instance locations helps prevent attribute leakage between overlapping bounding boxes and maintains structural consistency with the instances in the scene depth maps. Consequently, we employ the SAM (sam) model to ascertain the exact position of each instance. For an instance $\bm{i}$, by utilizing our generated scene depth map $\mathbf{m}_{scene}$ alongside its corresponding bounding box $\bm{b}_{i}$, we can segment the specific shape mask $\mathbf{m}_{i}$ of this instance, thereby facilitating subsequent merging:
@@
Merging rendering results. We employ the precise mask $\mathbf{m}_{i}$ obtained from SAM to constrain the rendering results of instance $\bm{i}$ to its own region, ensuring no influence on other instances. Specifically, we construct a new mask $\mathbf{m}_{i}^{\prime}$ by assigning a value of $\alpha$ to the areas where $\mathbf{m}_{i}$ equals $1$, while setting all other regions to -$\infty$. Simultaneously, we assign a background value of $\beta$ to the global rendering $\mathbf{r}_{c}$ through a mask $\mathbf{m}_{c}^{\prime}$. By applying the softmax function to the set $\{\mathbf{m}_{1}^{\prime},\mathbf{m}_{2}^{\prime},\ldots,\mathbf{m}_{n}^{\prime},\mathbf{m}_{c}^{\prime}\}$, we derive the spatial weights $\{\mathbf{m}_{1}^{\prime\prime},\mathbf{m}_{2}^{\prime\prime},\ldots,\mathbf{m}_{n}^{\prime\prime},\mathbf{m}_{c}^{\prime\prime}\}$ for each rendering instance. At each Cross Attention layer, the output can be expressed as follows to render the whole scene:
@@ 4.1 Implement Details
Tuning of text-to-depth models. We utilized a training set comprising 5,878 images from the LAION-art dataset (schuhmann2021laion), selecting only those with a resolution exceeding 512x512 pixels and an aesthetic score of $\geq{8.0}$. Depth maps for each image were generated using Depth Anything V2 (yang2024depthv2). Given the substantial noise present in the text descriptions associated with the images in LAION-art, we chose to produce corresponding image captions using BLIP2 (li2023blip2). We employed pyramid noise (pyramidnoise) to fine-tune the LDM3D model for 2,000 steps, utilizing the AdamW (adam) optimizer with a constant learning rate of $1e^{-4}$, a weight decay of $1e^{-2}$, and a batch size of 320.
Training of the layout-to-depth adapter. We adopted the MIGC (migc) architecture as the adapter for layout control. In alignment with this approach, we utilized the COCO dataset (coco) for training. We employed Stanza (stanza) to extract each instance description from the corresponding text for every image and used Grounding-DINO (gdino) to obtain the image layout. Furthermore, we augmented each instance’s description by incorporating modified versions that omitted adjectives, allowing our layout-to-depth adapter to focus more on global scene construction and the coarse-grained categories and structural properties of instances. We maintain the same batch size, learning rate, and other parameters as the previous work.
@@ 4.2 Experiment Setup
Baselines. We compared our proposed 3DIS method with state-of-the-art Multi-Instance Generation approaches. The methods involved in the comparison include training-free methods: BoxDiffusion (boxdiff) and MultiDiffusion (multidiffusion); and adapter-based methods: GLIGEN (gligen), InstanceDiffusion (instancediffusion), and MIGC (migc).
Evaluation Benchmarks. We conducted experiments using two widely adopted benchmarks, COCO-position (coco) and COCO-MIG (migc), to assess the performance of models in different aspects of instance generation. The COCO-position benchmark emphasizes the evaluation of a model’s capacity to control the spatial arrangement of instances, as well as their high-level categorical attributes. In contrast, the COCO-MIG benchmark is designed to test a model’s ability to precisely render fine-grained attributes for each generated instance. To rigorously compare the models’ performance in handling complex scene layouts, we concentrated our analysis on the COCO-position benchmark, specifically focusing on layouts containing five or more instances. For a comprehensive evaluation, each model generated 750 images across both benchmarks.
@@ 5 Conclusion
6 Appendix
6.1 Comparison of LDM3D and RichDreamer
Figure 9: Comparison of LDM3D and RichDreamer.
Upon investigation, we identified RichDreamer (richdreamer) and LDM3D (stan2023ldm3d) as the primary models employed for text-to-depth generation. To compare their performance, we utilized prompts from the COCO2014 dataset as input for both models, with the corresponding results illustrated in Fig. 9. Our analysis indicates that LDM3D demonstrates a superior ability to preserve the original SD1.4 priors, resulting in enhanced text comprehension and more precise control over scene generation. In contrast, RichDreamer exhibits certain shortcomings: (i) it often misses semantic details or omits entire objects in the depth maps, as seen in cases where essential elements like the cot and man are entirely absent; (ii) the depth maps produced by RichDreamer frequently suffer from artifacts such as blotches or thread-like distortions, particularly when used in conjunction with ControlNet. Therefore, after a thorough comparison, we selected LDM3D as the base model for text-to-depth generation in our 3DIS system.
6.2 Visualization on the impact of the LDM3D Fine-Tuning
Figure 10: Comparison of original LDM3D and finetuned LDM3D.
Although LDM3D is capable of generating relatively good depth maps, several issues remain: (i) Since LDM3D was trained using depth maps extracted from the DPT-Large Model (midas) , the resulting image quality is relatively poor. (ii) As a diffusion model trained by Gaussian noise, LDM3D exhibits limited ability to recover low-frequency content (offsetnoise). This is clearly illustrated in Fig. 10, where the generated depth maps struggle to produce large uniform color blocks. Moreover, the average color value of the depth maps tends to converge towards the initial noise, whose mean value is close to 0. This constraint places a harmful limitation on text-to-depth generation.
To address (i), we fine-tuned the model using depth maps extracted from the latest Depth-Anything V2 model. For (ii), we adopted pyramid noise instead of Gaussian noise, which helps mitigate the constraints on text-to-depth generation. As shown in Fig. 10, the fine-tuned LDM3D model is capable of generating depth maps with higher contrast and improved overall quality.
6.3 More results of the generated scene depth map
Figure 11: More results of the generated scene depth map.
Fig. 11 presents additional results of scene depth map generation using our 3DIS system. The results demonstrate that, even with complex layouts, 3DIS effectively understands and generates cohesive scenes, harmoniously placing all objects within them. Furthermore, even in cases of significant overlap, such as the five suitcases in the fifth row, 3DIS handles the arrangement with precision, maintaining clear object separation and preventing blending.
6.4 More results on COCO-MIG benchmark
Fig. 12 presents additional results of 3DIS on the COCO-MIG dataset, revealing several key advantages over the previous state-of-the-art model, MIGC. First, 3DIS demonstrates superior scene construction capabilities, as seen in the first and second rows, where it constructs more coherent scenes that appropriately place all specified instances—such as rendering an indoor environment when prompted with “refrigerator.” Second, 3DIS exhibits enhanced detail rendering, as shown in the fourth to sixth rows. By leveraging the more advanced SDXL model in a training-free manner, 3DIS outperforms MIGC, which primarily relies on SD1.5, producing more visually appealing and structurally refined results. Third, 3DIS handles smaller instances better, as demonstrated in the third row with the “red bird” and “yellow dog.” Its ability to render at higher resolutions using SDXL leads to clearer and more accurate depictions of these smaller objects. Finally, 3DIS excels in managing overlapping objects, as illustrated in the seventh row, where it avoids object merging while generating the scene’s depth map.
Figure 12: More qualitative results on the COCO-MIG.
6.5 Limitation
Although 3DIS leverages various foundation models for rendering fine instance details, its scene construction continues to rely on the less advanced SD1.5 model. This dependency limits 3DIS’s capacity to accurately generate complex structures, particularly those that SD1.5 struggles with, such as intricate shapes or highly detailed spatial configurations. As a result, 3DIS may occasionally fail in scenarios that demand more sophisticated generative capabilities. Addressing this limitation in future work could involve the development of specialized datasets aimed at enhancing the model’s proficiency in handling complex geometries, thereby improving the overall robustness and versatility of 3DIS in a broader range of applications.